# PANORAMIC: Discovery of an Ultra-Massive Grand-Design Spiral Galaxy at

### v1

@@
We report the discovery of an ultra-massive grand-design red spiral galaxy, named Zhúlóng (Torch Dragon), at in the JWST PANORAMIC survey, identified as the most distant bulge+disk galaxy with spiral arms known to date. Zhúlóng displays an extraordinary combination of properties: 1) a classical bulge centered in a large, face-on exponential stellar disk (half-light radius of ), with spiral arms extending across 19 kpc; 2) a clear transition from the red, quiescent core ( mag) with high stellar mass surface density () to the star-forming outer regions, as revealed by spatially resolved SED analysis, which indicates significant inside-out galaxy growth; 3) an extremely high stellar mass at its redshift, with comparable to the Milky Way, and an implied baryon-to-star conversion efficiency () that is 1.5 times higher than even the most efficient galaxies at later epochs; 4) despite an active disk, a relatively modest overall star formation rate (), which is 0.5 dex below the star formation main sequence at and 10 times lower than ultra-massive dusty galaxies at . Altogether, Zhúlóng shows that mature galaxies emerged much earlier than expected in the first billion years after the Big Bang through rapid galaxy formation and morphological evolution. Our finding offers key constraints for models of massive galaxy formation and the origin of spiral structures in the early universe.
@@
galaxies: high-redshift – galaxies: evolution – galaxies: star-formation – galaxies: photometry1 Introduction
@@
At the forefront of these discoveries are massive galaxies. Panchromatic JWST imaging and spectroscopy from the observed optical to mid-infrared have confirmed not only that ultra-massive galaxies existed in the early Universe ( at ), but that they may also exist at unexpectedly high abundance (e.g. Labbé et al., 2023; Xiao et al., 2024; Weibel et al., 2024b). They are so massive, with almost Milky Way (MW) mass only at 1 Gyr after the Big Bang, suggesting that galaxies are forming in a more efficient way than previously thought. Based on their existence in small survey areas (typical JWST imaging campaigns from Cycle 1 sq arcmin per field), these ultra-massive galaxies, like the ones reported in Xiao et al. (2024), require on average about 50% of the baryons in their dark matter halos to be converted into stars – two to three times more than even the most efficient galaxies at later times (Moster et al., 2013; Wechsler & Tinker, 2018; Shuntov et al., 2022). While questions remain about incomplete data or modeling assumptions for photometric samples (e.g. Wang et al., 2024a; Williams et al., 2024a; Desprez et al., 2024), cosmic variance from small fields and sample sizes also complicate interpretation. This motivates studies of known massive candidates, because if true, these findings definitively raise new important questions about the theory of evolutionary drivers of early galaxies.
@@
Here, we report on the serendipitous discovery of a unique ultra-massive galaxy at in the PANORAMIC111Parallel wide-Area Nircam Observations to Reveal And Measure the Invisible Cosmos (PANORAMIC) Survey (Williams et al., 2024b). Unlike the other ultra-massive candidates described above, it has a striking evolved morphology: a quiescent-like classical bulge + star-forming stellar disk + grand-design spiral arms (defined as spiral structures spanning the whole galaxy, from the nucleus to outskirts, with spiral arms starting at diametrically opposite points; e.g. Binney & Tremaine 2008; Dobbs & Baba 2014; Sellwood & Masters 2022), already at 1 billion years after the Big Bang. More interestingly, it is very large in stellar size (19 kpc in diameter), extremely massive (, indicating high baryon-to-star formation efficiency), with a modest star formation rate with red color (; F150WF444W mag). Considering only a few stellar spiral galaxies have been identified at (Jain & Wadadekar, 2024; Wang et al., 2024b; Umehata et al., 2024; Costantin et al., 2023; Wu et al., 2023), this source becomes the most distant massive spiral galaxies discovered so far, providing direct empirical evidence of the rapid dynamical evolution of massive galaxies at . Given its unique red color, morphology, and stellar properties, we name it Zhúlóng222Zhúlóng ({CJK}UTF8gkai烛龙, or Torch Dragon), which is a giant red solar dragon and god in Chinese mythology. It supposedly had a human’s face and snake’s body, created day and night by opening and closing its eyes, and created seasonal winds by breathing..
@@ 2 Observational data
Zhúlóng (, [J2000] = 150.124874, 2.092919) was serendipitously discovered in the field (association) named j100024p0208 in the recent data release of the JWST PANORAMIC survey (GO-2514; PIs: C. Williams & P. Oesch; Williams et al., 2024b), a pure parallel extragalactic NIRCam imaging program. Zhúlóng further lies within NIRCam imaging footprint of the COSMOS-Web field (Casey et al., 2023), but is not covered by the COSMOS-PRIMER program (Dunlop et al., 2021).
@@ 2.1 HST and JWST NIRCam observations
Multi-wavelength photometric measurements are derived following the same procedure as outlined in Weibel et al. (2024b). Briefly, we use SExtractor (Bertin & Arnouts, 1996) in dual image mode with an inverse-variance weighted stack of the longest wavelength NIRCam wide filters, F277W+F356W+F444W, as the detection image. In this study, fluxes are measured in 016, 035, 05, and 07 radius circular apertures in images that are point spread function (PSF)-matched to the F444W band. Total fluxes are derived from the Kron AUTO aperture provided by SExtractor in the F444W band, in addition to a correction based on the encircled energy of the Kron aperture on the F444W PSF. Detailed descriptions of data reduction and photometric measurements are provided in Williams et al. (2024b).
@@ 2.2 Longer-wavelength observations
Although Zhúlóng lies within the COSMOS-Web field, it falls in a gap in the MIRI F770W footprints. Observations at longer wavelengths cover the location of Zhúlóng, but none return a detection, which include Spitzer MIPS 24 µm (Le Floc’h et al., 2009), Herschel PACS (Lutz et al., 2011) and SPIRE (Oliver et al., 2012), NIKA2 Cosmological Legacy Survey at 1.2mm and 2.0mm (rms of 0.3 and 0.09 mJy beam-1, respectively; Bing et al., 2023, Carvajal Bohorquez et al. in prep), ALMA 2.0mm (rms of 0.12 mJy beam-1; Project ID: 2021.1.00705.S), ALMA 3.0mm (rms of 0.06 mJy beam-1; Project ID: 2021.1.01005.S), and VLA 3 GHz (Smolčić et al., 2017) and 1.4 GHz (Schinnerer et al., 2010). More recently, the ALMA large program CHAMPS (PI: A. Faisst; priv. comm.; Faisst et al. in prep) also covers Zhúlóng at 1.2 mm, but again it remains undetected. The local rms level is 0.2 mJy beam-1 in the 1.2mm map of spatial resolution. The 3 upper limit at 1.2mm is 0.6 mJy beam-1, assuming the dust distribution is unresolved, indicating that Zhúlóng has low far-infrared luminosity.
@@ 2.3 Zhúlóng with robust photometric redshift:
The photometric redshift of Zhúlóng is estimated using the SED-fitting code EAZY (Brammer et al., 2008), with the blue_sfhz_13 template set which imposes redshift-dependent star formation histories (SFHs), excluding SFHs that start earlier than the age of the Universe at a given redshift333https://github.com/gbrammer/eazy-photoz/tree/master/templates/sfhz. We apply an error floor of 5% prior to running eazy to account for possible remaining systematic uncertainties in the photometry and to allow for more flexibility in the SED-fitting. The redshift range for fitting is set to . Additional details on the photometric redshift modeling are presented in Williams et al. (2024b).
The EAZY best-fit results are shown in Fig. 1. The redshift probability distribution for Zhúlóng is strongly constrained by a prominent Balmer/4000 Å break observed in the central region (0.16′′ radius), resulting in narrow uncertainties with a single redshift solution . In addition, the neighboring south-west clump has evidence of strong emission lines, which boost the photometry in the F277W and F444W bands, leading to an even better constrained photometric redshift: . This consistency illustrates the reliability of Zhúlóng. Nevertheless, it will be important to confirm the redshift through spectroscopy in the future, such as with the upcoming COSMOS-3D grism observations (GO-5893, Kakiichi et al. 2024) or with NIRSpec spectra.
@@ 3.1 Appearance of bulge, disk, and spiral structures
Fig. 2 shows the best-fit results of the stacked image. From the residual maps, Zhúlóng shows evidence for a grand-design spiral arm structure, with high contrast primary arms extending from the bulge pointing north and south (e.g. Dobbs & Baba, 2014; Sellwood & Masters, 2022). The presence of the spiral arms is also indicated in the surface brightness profiles, which show the discrepancy between the observed points and the best-fit model. The single Sérsic fit yields kpc, suggesting that Zhúlóng has a large stellar size. The large reveals that the stellar light is more concentrated at the galaxy center, indicating the potential existence of a bulge structure. Meanwhile, the significant inhomogeneity in the core of the residual map also suggests the need for an additional bulge component in the fitting process, which is mitigated in the residual map of the double Sérsic fit. On the other hand, the double Sérsic fit yields evidence for both bulge + disk structures, with spiral arms shown in the residual map. The bulge component has , close to the shape of classical bulges in the local Universe (), suggesting the existence of a well-evolved bulge even in the first Gyr after the Big Bang. The bulge-to-total (light) ratio is . The disk component has , kpc, and , indicating the existence of a large face-on exponential stellar disk. In addition, we note that both of the residual maps show clear evidence of stellar clumps along the spiral arms, as observed in galaxies at cosmic noon (e.g., Rujopakarn et al., 2016; Claeyssens et al., 2023, 2024) and as also found in numerical simulations (e.g., Tamburello et al., 2015; Mandelker et al., 2017; Fensch & Bournaud, 2021; Ceverino et al., 2023; Renaud et al., 2024). Furthermore, we also perform the morphological analysis for individual images, and the results remain the same (see Appendix A).
@@
Given the extended nature of Zhúlóng, and the amount of flux present in spiral arms that are not well-accounted for by either Sérsic model, we also measure the size that encloses 80% of the light empirically, using the azimuthally-averaged light profile from the stacked image without any assumption on the parametric shape of the light distribution. Using this method, we find that these radii are larger than inferred assuming the parametric Sérsic (or double Sérsic) model. We measure R80=9.5 kpc for the radii enclosing 80% of the light. R80 is extremely large compared to its half-light radius (while better reflecting the overall extent of stellar light) reflecting the fact that the spatial extent of the stars in the galaxy reaches more than 19 kpc in diameter.
@@ 3.2 A quiescent-like bulge and star-forming disk
In Fig. 3, from the central core to the outer region, the best-fit SED changes significantly from quiescent to star-forming populations. The core is dominated by old stellar populations with a red color ( ABmag). The best-fit SED exhibits a strong 4000 Å break, very low (sSFR/yr, and a lack of emission lines. The observed colors are consistent with NIRCam-based photometric selections for quiescent galaxies (Long et al., 2024). The rest-frame U-V and V-J colors as measured from eazy are 1.4 and 1.0 mag, respectively, consistent with a quiescent/post-starburst fading stellar population (Williams et al., 2009). The restframe colors are also consistent with extensions of UVJ color criteria for quiescent galaxies that have been adapted to to accommodate younger passively evolving galaxies (Antwi-Danso et al., 2023). In contrast, the disk measured in an annulus between 0.5 is dominated by younger stellar populations with a relatively bluer color ( ABmag) and, unlike the inner region, allows the presence of weak emission lines in the SED.
@@ 3.3 Ultra-massive spiral galaxy with accelerated formation
Here we focus on the integrated properties of Zhúlóng. The total , , and SFR are derived by fitting the UV-to-FIR SED from the JWST+HST+ALMA photometry, using Bagpipes. Similar to Sect. 3.2, we use the Calzetti et al. (2000) reddening law, the Bruzual & Charlot (2003) stellar population models, and the Draine & Li (2007) dust emission model (with energy balance), assuming different SFH (double-power-law SFH and delayed SFH) and free metallicity. The derived main properties with delayed SFH are listed in Table 1. To test the result, we also perform SED fitting with CIGALE (Boquien et al., 2019), which produces very consistent values within the errors. The total stellar mass is also consistent with the sum of the individual masses obtained from the different annuli (see Sect. 3.2). In general, we find that Zhúlóng is extremely massive with .
@@
In addition, in Fig. 4, we also collect examples from the literature of galaxies whose stellar structures are consistent with spiral arms at , the highest redshift at which spiral structure has been detected in stars (rather than just gas) to date (Jain & Wadadekar, 2024; Wang et al., 2024b; Umehata et al., 2024; Costantin et al., 2023; Wu et al., 2023). We note that some other galaxies are reported to show potential spiral structures in JWST images, but are not shown in our figure due to the lack of comparable analyses and stellar mass values reported (e.g., Ferreira et al., 2023). In addition, sources showing only stellar disks but without clear spiral arms are also not in our plot (e.g., Yan et al., 2024). Compared with the literature-reported spiral galaxies, Zhúlóng appears to be the most distant spiral galaxy discovered so far.

### v2

@@
We report the discovery of an ultra-massive grand-design red spiral galaxy, named Zhúlóng (Torch Dragon), at in the JWST PANORAMIC survey, identified as the most distant bulge+disk galaxy candidate with spiral arms known to date. Zhúlóng displays an extraordinary combination of properties: 1) a classical bulge centered in a large, face-on exponential stellar disk (half-light radius of ), with spiral arms extending across 19 kpc; 2) a clear transition from the red, quiescent core ( mag) with high stellar mass surface density () to the star-forming outer regions, as revealed by spatially resolved SED analysis, which indicates significant inside-out galaxy growth; 3) an extremely high stellar mass at its redshift, with comparable to the Milky Way, and an implied baryon-to-star conversion efficiency () that is 1.5 times higher than even the most efficient galaxies at later epochs; 4) despite an active disk, a relatively modest overall star formation rate (), which is 0.5 dex below the star formation main sequence at and 10 times lower than ultra-massive dusty galaxies at . Altogether, Zhúlóng shows that mature galaxies emerged much earlier than expected in the first billion years after the Big Bang through rapid galaxy formation and morphological evolution. Our finding offers key constraints for models of massive galaxy formation and the origin of spiral structures in the early universe.
@@
galaxies: high-redshift – galaxies: formation – galaxies: star-formation – galaxies: spiral – galaxies: photometry1 Introduction
@@
At the forefront of these discoveries are massive galaxies. Panchromatic JWST imaging and spectroscopy from the observed optical to mid-infrared have confirmed not only that ultra-massive galaxies existed in the early Universe ( at ), but that they may also exist at unexpectedly high abundance (e.g. Labbé et al., 2023; Xiao et al., 2024; Weibel et al., 2024b). They are so massive, with almost Milky Way (MW) mass only at 1 Gyr after the Big Bang, suggesting that galaxies are forming in a more efficient way than previously thought. Based on their existence in small survey areas (typical JWST imaging campaigns from Cycle 1 sq arcmin per field), these ultra-massive galaxies, like the ones reported in Xiao et al. (2024), require on average about 50% of the baryons in their dark matter halos to be converted into stars – two to three times more than even the most efficient galaxies at later times (Moster et al., 2013; Wechsler & Tinker, 2018; Shuntov et al., 2022). While questions remain about incomplete data or modeling assumptions for photometric samples (e.g. Wang et al., 2024a; Williams et al., 2024; Desprez et al., 2024), cosmic variance from small fields and sample sizes also complicate interpretation. This motivates studies of known massive candidates, because if true, these findings definitively raise new important questions about the theory of evolutionary drivers of early galaxies.
@@
Here, we report on the serendipitous discovery of a unique ultra-massive galaxy at in the PANORAMIC111Parallel wide-Area Nircam Observations to Reveal And Measure the Invisible Cosmos (PANORAMIC) Survey (Williams et al., 2025). Unlike the other ultra-massive candidates described above, it has a striking evolved morphology: a quiescent-like classical bulge + star-forming stellar disk + grand-design spiral arms (defined as spiral structures spanning the whole galaxy, from the nucleus to outskirts, with spiral arms starting at diametrically opposite points; e.g. Binney & Tremaine 2008; Dobbs & Baba 2014; Sellwood & Masters 2022), already at 1 billion years after the Big Bang. More interestingly, it is very large in stellar size (19 kpc in diameter), extremely massive (, indicating high baryon-to-star formation efficiency), with a modest star formation rate with red color (; F150WF444W mag). Considering only a few stellar spiral galaxies have been identified at (Jain & Wadadekar, 2024; Wang et al., 2024b; Umehata et al., 2024; Costantin et al., 2023; Wu et al., 2023), this source becomes the most distant massive spiral galaxies discovered so far, providing direct empirical evidence of the rapid dynamical evolution of massive galaxies at . Given its unique red color, morphology, and stellar properties, we name it Zhúlóng222Zhúlóng ({CJK}UTF8gkai烛龙, or Torch Dragon), which is a giant red solar dragon and god in Chinese mythology. It supposedly had a human’s face and snake’s body, created day and night by opening and closing its eyes, and created seasonal winds by breathing..
@@ 2 Observational data
Zhúlóng (, [J2000] = 150.124874, 2.092919) was serendipitously discovered in the field (association) named j100024p0208 in the recent data release of the JWST PANORAMIC survey (GO-2514; PIs: C. Williams & P. Oesch; Williams et al., 2025), a pure parallel extragalactic NIRCam imaging program. The source was originally identified by chance during a search for massive galaxy candidates that were color-selected to exhibit a strong Balmer break at (Long et al., 2024). Its spiral structure was identified during visual inspection of those candidates, and was unique among the sample. Zhúlóng lies within the NIRCam imaging footprint of the COSMOS-Web field (Casey et al., 2023), but is not covered by the COSMOS-PRIMER program (Dunlop et al., 2021).
@@ 2.1 HST and JWST NIRCam observations
Multi-wavelength photometric measurements are derived following the same procedure as outlined in Weibel et al. (2024b). Briefly, we use SExtractor (Bertin & Arnouts, 1996) in dual image mode with an inverse-variance weighted stack of the longest wavelength NIRCam wide filters, F277W+F356W+F444W, as the detection image. In this study, fluxes are measured in 016, 035, 05, and 07 radius circular apertures in images that are point spread function (PSF)-matched to the F444W band. Total fluxes are derived from the Kron AUTO aperture provided by SExtractor in the F444W band, in addition to a correction based on the encircled energy of the Kron aperture on the F444W PSF. Detailed descriptions of data reduction and photometric measurements are provided in Williams et al. (2025).
@@ 2.2 Longer-wavelength observations
Although Zhúlóng lies within the COSMOS-Web field, it falls in a gap in the MIRI F770W footprints. Observations at longer wavelengths cover the location of Zhúlóng, but none return a detection, which include Spitzer MIPS 24 µm (Le Floc’h et al., 2009), Herschel PACS (Lutz et al., 2011) and SPIRE (Oliver et al., 2012), NIKA2 Cosmological Legacy Survey at 1.2mm and 2.0mm (rms of 0.3 and 0.09 mJy beam-1, respectively; Bing et al., 2023, Carvajal Bohorquez et al. in prep), ALMA 2.0mm (rms of 0.12 mJy beam-1; Project ID: 2021.1.00705.S), ALMA 3.0mm (rms of 0.06 mJy beam-1; Project ID: 2021.1.01005.S), and VLA 3 GHz (Smolčić et al., 2017) and 1.4 GHz (Schinnerer et al., 2010). More recently, the ALMA large program CHAMPS (PI: A. Faisst; priv. comm.; Faisst et al. in prep) also covers Zhúlóng at 1.2 mm, but again it remains undetected. The local rms level is 0.2 mJy beam-1 in the 1.2mm map of spatial resolution. The 3 upper limit at 1.2mm is 0.6 mJy, assuming the dust distribution is unresolved, indicating that Zhúlóng has low far-infrared luminosity.
@@ 2.3 Zhúlóng with robust photometric redshift:
The photometric redshift of Zhúlóng is estimated using the SED-fitting code EAZY (Brammer et al., 2008), with the blue_sfhz_13 template set which imposes redshift-dependent star formation histories (SFHs), excluding SFHs that start earlier than the age of the Universe at a given redshift333https://github.com/gbrammer/eazy-photoz/tree/master/templates/sfhz. We apply an error floor of 5% prior to running eazy to account for possible remaining systematic uncertainties in the photometry and to allow for more flexibility in the SED-fitting. The redshift range for fitting is set to . Additional details on the photometric redshift modeling are presented in Williams et al. (2025).
The EAZY best-fit results are shown in Fig. 1. The redshift probability distribution for Zhúlóng is strongly constrained by a prominent Balmer/4000 Å break observed in the central region (0.16′′ radius), resulting in narrow uncertainties with a single redshift solution . In addition, the neighboring south-west clump has evidence of strong emission lines, which boost the photometry in the F277W and F444W bands, leading to an even better constrained photometric redshift: . This consistency illustrates the reliability of Zhúlóng. Further tests and a detailed discussion of the robustness of the redshift are presented in Appendix A. Nevertheless, we underline that this is still a photometric redshift. It will be important to confirm the redshift through spectroscopy in the future, such as with the upcoming COSMOS-3D grism observations (GO-5893, Kakiichi et al. 2024) or with NIRSpec spectra.
@@ 3.1 Appearance of bulge, disk, and spiral structures
Fig. 2 shows the best-fit results of the stacked image. From the residual maps, Zhúlóng shows evidence for a grand-design spiral arm structure, with high contrast primary arms extending from the bulge pointing north and south (e.g. Dobbs & Baba, 2014; Sellwood & Masters, 2022). The presence of the spiral arms is also indicated in the surface brightness profiles, which show the discrepancy between the observed points and the best-fit model. The single Sérsic fit yields kpc, suggesting that Zhúlóng has a large stellar size. The large reveals that the stellar light is more concentrated at the galaxy center, indicating the potential existence of a bulge structure. Meanwhile, the significant inhomogeneity in the core of the residual map also suggests the need for an additional bulge component in the fitting process, which is mitigated in the residual map of the double Sérsic fit. On the other hand, the double Sérsic fit yields evidence for both bulge + disk structures, with spiral arms shown in the residual map. The bulge component has , close to the shape of classical bulges in the local Universe (), suggesting the existence of a well-evolved bulge even in the first Gyr after the Big Bang. The bulge-to-total (light) ratio is . The disk component has , kpc, and , indicating the existence of a large face-on exponential stellar disk. In addition, we note that both of the residual maps show clear evidence of stellar clumps along the spiral arms, as observed in galaxies at cosmic noon (e.g., Rujopakarn et al., 2016; Claeyssens et al., 2023, 2024) and as also found in numerical simulations (e.g., Tamburello et al., 2015; Mandelker et al., 2017; Fensch & Bournaud, 2021; Ceverino et al., 2023; Renaud et al., 2024). Furthermore, we also perform the morphological analysis for individual images, and the results remain the same (see Appendix B).
@@
Given the extended nature of Zhúlóng, and the amount of flux present in spiral arms that are not well-accounted for by either Sérsic model, we also measure the size that encloses 80% of the light empirically, using the azimuthally-averaged light profile from the stacked image without any assumption on the parametric shape of the light distribution. Using this method, we find that these radii are larger than inferred assuming the parametric Sérsic (or double Sérsic) model. We measure R80=9.5 kpc for the radii enclosing 80% of the light. R80 is extremely large compared to its half-light radius (while better reflecting the overall extent of stellar light), reflecting the fact that the spatial extent of the stars in the galaxy reaches more than 19 kpc in diameter.
@@ 3.2 A quiescent-like bulge and star-forming disk
In Fig. 3, from the central core to the outer region, the best-fit SED changes significantly from quiescent to star-forming populations. The core is dominated by old stellar populations with a red color ( ABmag). The best-fit SED exhibits a strong Balmer/4000 Å break, very low (sSFR/yr, and a lack of emission lines. The observed colors are consistent with NIRCam-based photometric selections for quiescent galaxies (Long et al., 2024). The rest-frame U-V and V-J colors as measured from eazy are 1.4 and 1.0 mag, respectively, consistent with a quiescent/post-starburst fading stellar population (Williams et al., 2009). The restframe colors are also consistent with extensions of UVJ color criteria for quiescent galaxies that have been adapted to to accommodate younger passively evolving galaxies (Antwi-Danso et al., 2023). In contrast, the disk measured in an annulus between 0.5 is dominated by younger stellar populations with a relatively bluer color ( ABmag) and, unlike the inner region, allows the presence of weak emission lines in the SED.
@@ 3.3 Ultra-massive spiral galaxy with accelerated formation
Here we focus on the integrated properties of Zhúlóng. The total , , and SFR are derived by fitting the UV-to-FIR SED from the JWST+HST+ALMA photometry, using Bagpipes. Similar to Sect. 3.2, we use the Calzetti et al. (2000) reddening law, the Bruzual & Charlot (2003) stellar population models, and the Draine & Li (2007) dust emission model (with energy balance), assuming different SFH (double-power-law SFH and delayed SFH) and free metallicity. The derived main properties with delayed SFH are listed in Table 1. To test the result, we also perform SED fitting with CIGALE (Boquien et al., 2019), which produces very consistent values within the errors. The total stellar mass is also consistent with the sum of the individual masses obtained from the different annuli (; see Sect. 3.2). In general, we find that Zhúlóng is extremely massive with .
@@
In addition, in Fig. 4, we compile all currently known examples from the literature of galaxies whose stellar structures are consistent with spiral arms at , the highest redshift at which spiral structure has been detected in stars (rather than just gas) to date (Jain & Wadadekar, 2024; Wang et al., 2024b; Umehata et al., 2024; Costantin et al., 2023; Wu et al., 2023). We note that while some other galaxies have been reported to show potential spiral-like structures in JWST images, they are not included in our figure due to the lack of comparable analyses and stellar mass estimates (e.g., Ferreira et al., 2023). Thus, we caution that there may be more spiral galaxies at than those currently identified, and a systematic study of high-redshift spiral galaxies is necessary to confirm this. In addition, sources showing only stellar disks but without clear spiral arms are also not in our plot (e.g., Yan et al., 2024). Compared with the literature-reported spiral galaxies, Zhúlóng appears to be the most distant spiral galaxy discovered so far.